# Bad Students Make Great Teachers: Active Learning Accelerates Large-Scale Visual Understanding

### v2

@@ Abstract
We propose a method for accelerating large-scale pre-training with online data selection policies. For the first time, we demonstrate that model-based data selection can reduce the total computation needed to reach the performance of models trained with uniform sampling. The key insight which enables this “compute-positive” regime is that small models provide good proxies for the loss of much larger models, such that computation spent on scoring data can be drastically scaled down but still significantly accelerate training of the learner. These data selection policies also strongly generalize across datasets and tasks, opening an avenue for further amortizing the overhead of data scoring by re-using off-the-shelf models and training sequences. Our methods, ClassAct and ActiveCLIP, require 46% and 51% fewer training updates and up to 25% less total computation when training visual classifiers on JFT and multimodal models on ALIGN, respectively. Finally, our paradigm seamlessly applies to the curation of large-scale image-text datasets, yielding a new state-of-the-art in several multimodal transfer tasks and pretraining regimes.
**footnotetext: Equal technical contribution. ${}^{\dagger}$urrent affiliation: Harvard University, work done while at Google DeepMind.
@@ 1 Introduction
In recent years, the explosion of available data has catalyzed the development of increasingly sophisticated models across domains. However, maintaining this progress comes at a cost—power-law scaling for vision and language models [19, 44] dictates that each increment in performance will require another order of magnitude in computation.
One of the key assumptions of these empirical power-laws is that training data is sampled uniformly throughout training. In contrast, active learning approaches sample data in a prioritized fashion, such that computation is spent training on the data that maximally contributes to task performance [22, 25, 35], with the ultimate goals of improving data efficiency and reducing the cost of training runs.
@@
Figure 1: Active learning accelerates large scale visual understanding. For large scale classification and multimodal learning tasks, prioritised training on data selected using our active selection methods ClassAct (top) and ActiveCLIP (bottom) requires significantly fewer updates than training on data sampled IID.
@@
In the first instance, data selection based on hand-engineered features, such as removing images that are the not of a specified shape or that only contain a single colour [2] can trivially improve training efficiency at minimal computational overhead. However, such heuristics are limited in their effectiveness by the expertise of the human designer and are not guaranteed to transfer to training of different models, data modalities or tasks.
@@
In-contrast, recent model-based filtering heuristics have shown promise in small-to-medium visual classification tasks by focusing training on ‘hard’ and omitting ‘easy’ data [36], as evaluated by the loss of the learner model itself. Similar gains in data efficiency have been observed in language modeling, where pre-trained models are employed to exclude low-quality data [12]. However, these methods often spend as much computation on the curation of datasets as is gained from subsequent pretraining, with the consequence that training is overall less compute-efficient than training on uniformly-sampled (IID) data.
In this work we investigate whether data-selection policies can instead be designed to be compute-positive, such that the sum of all computation needed to reach a target performance, including the pre-training of reference models and inference required for data selection, is less than the computation incurred by an equivalent IID training run.
To address this question, we explore a new model-based active learning framework in which data is scored for prioritised sampling using models that are significantly smaller and cheaper than the learner model being trained. The contributions of our work are summarized below:
  1. 1.
@@
    Benchmarking heuristics for large-scale pretraining: We conduct an initial investigation of various scoring functions that quantify the importance of examples during training, including loss-based prioritization [20, 17, 13, 36] and ‘learnability scores’ [28]. From this we find that noise-removal with pretrained models to be an essential component for accelerating learning from large-scale datasets, with efficiency gains up to 46%.
  2. 2.
    Generalizing selection policies across model scale: We investigate whether scoring metrics derived from a particular model scale generalize to larger models, and find learnability scores to be uniquely robust, with 1000$\times$ smaller scoring models still providing significant efficiency gains. Thanks to these efficient scoring models, we reach the compute-positive regime, matching the baseline performance with 25% fewer FLOPs.
  3. 3.
@@
    Multimodal data curation: We find our framework to provide a powerful tool for multimodal data curation. Using reference models pretrained on small, clean datasets, we substantially accelerate and improve pretraining on much larger, noisier datasets, yielding strong performance on several multimodal understanding tasks.
  4. 4.
@@
    Amortizing data selection policies: We demonstrate that data-selection policies can be trained on one task, but used to accelerate the training of subsequent models on different but related tasks, suggesting that such policies can be easily derived from pre-trained models.
  5. 5.
@@
    Simplifying active-learning with reference models: Lastly, we demonstrate that pre-trained reference models may not be necessary at all, where these models are small and can be trained in parallel in an accelerated manner on larger batches than the learner model, while still being compute-positive.
@@ 2 Related Work
Data pruning. One approach to data-selection is to identify and sub-select data ahead of training. Paul et al. [29] show that norms of the gradient or prediction error can be used to discard around half of CIFAR10 with no loss in performance. Sorscher et al. [36] go further and argue that data pruning to create ‘foundation datasets’ can in theory reduce neural scaling laws from power-law to exponential given access to a high-quality scoring metric. These methods have since been deployed for the curation of web-scale datasets in both language modeling [27] and multimodal learning [1, 26], demonstrating large reductions in the amount of data required together with performance improvements. A potential criticism of this line of work is the repeated observation of sequential concept learning even in an IID training run [31, 3, 7], which static data-filtering pipelines cannot accommodate due to their inability to adapt to the state of the model. More critically, in the single-epoch regime that is becoming typical of large model training [14, 19], pre-filtering can be as expensive as learning from it, a shortcoming which we address in this work.
Online active learning. The natural alternative to the ‘foundation dataset’ approach is to continuously filter data throughout training. Unlike data-pruning, online active learning applies naturally to the semi-infinite / single-epoch regime. One approach (Online Batch Selection [23]) is to score, filter and train all using the same learner model, which has the theoretical advantage that the importance of data can be determined relative to the current state of the learner. In terms of metrics, the Reducible Holdout Loss (RHO) [28] also uses the concept of a reference model to identify learnable data points not yet well represented. Other proposed heuristics include memorization for long-tailed data [10] and assigning “complexity” scores based on the number of times the example is forgotten during training [40]. None of these approaches however have demonstrated that the cost of scoring can be mitigated to the point of justifying learner efficiency gains. Most related to our work is DoReMi [41], which uses a small proxy model in order to determine the optimal data-mixture weights for the subsequent training of a larger model.
Meta-learning. Finally, there have also been combinations of related ideas in the meta-learning literature, which share the goal of improving overall training efficiency at the cost of per-update complexity. An example is Tack et al. [39], who introduce online data selection at each step within a short optimisation horizon and demonstrate both impressive memory savings and improved performance. For Q&A tasks in the LLM space, Hu et al. [15] show how a meta-learned weighting model can learn to mask the contribution of unimportant sub-sequences to the overall loss, improving results.
@@ 3 Methods
In this section we discuss the motivation, design and practical implementation of our active learning framework. In Section 3.1, we discuss how to parallelise online data selection alongside training. Section 3.2 examines which statistics could be relevant for large-scale data selection. Section 3.3 analyses the cost of existing AL frameworks and motivates the introduction of our own. Section 3.4 describes the details relating to applying our framework to large-scale classification and multi-modal learning.
@@ 3.1 Data selection as prioritized replay
We approach this problem from the reinforcement learning perspective, where data generation (‘acting’) and learning processes are decoupled.
We apply this framework to standard visual learning tasks in the following way: we sample uniformly from the training set ${\bm{x}}_{i}\sim\mathcal{D}$ and assign a score $s_{i}=s({\bm{x}}_{i}|\theta)$ to each data point ${\bm{x}}_{i}$ using model parameters $\theta$. Given a large enough collection of scored examples stored in a memory bank =$\mathcal{M}=\{{\bm{x}}_{i}\}_{i\in(0,\dots,M-1)}$, we sample non-uniformly according to their scores [32]. Specifically, we sample ${\bm{x}}_{i}\stackrel{{\scriptstyle\pi}}{{\sim}}\mathcal{M}$, where the policy $\pi$ is computed as a softmax over scores:
$\pi({\bm{x}}_{i})=\frac{\exp(s_{i}({\bm{x}}_{i}))}{\sum_{j}\exp(s_{j}({\bm{x}}_{j}))}$ (1)
The ‘actors’ continuously read data, compute scores, and write the scored data to the memory bank in parallel to the learning process. This allows maximum utilization of the hardware accelerators needed for learning by offloading the data processing and scoring tasks to actors which can utilize more cost-effective asynchronous computation.
@@ 3.2 Statistics for data selection
can use the current parameters of the learner $\theta^{t}$ or those of a fixed model $\theta^{*}$.
The opposite argument can been made for favoring examples that are easily solved by a well-trained model, as such a prioritization removes the noisy examples present in large-scale datasets:
@@
This prioritization scheme is commonly used in multimodal learning for identifying high-quality examples with pre-trained models [13, 33, 34].
@@
A special case of learnability scores (the Reducible Hold-Out Loss [28]) uses a model $\theta^{\textrm{ho}}$ specifically trained on a held-out dataset to ensure the independence of its predictions from those of the current learner $s^{\textrm{learn}}({\bm{x}}_{i}|\theta^{t},\theta^{\textrm{ho}})$. We assess in Section 4.2 whether this is necessary when training on large-scale image datasets.
Figure 2: Amortizing the cost of data selection. Draw to scale - length of bars proportional to FLOP counts from Figure 4. Expensive model policies (e.g. using a ViT-B to score data for the ViT-L learner, or ‘B normal-$\rightarrow$ L’) produce better learner speedups, at the expense of the additional FLOPs associate with data selection. However, some or part of these costs can be amortized, depending on the context. Using off-the-shelf reference models removes the need to train from scratch (yellow). Since the reference model is fixed throughout training, scores can be assigned once to a ‘foundation dataset’ and amortized across many training runs (lime green; [36, 28]). Since the online model is independent of the learner model and generalizes across scale, data selection policies can also be distilled as a fixed ordering of a given dataset (a ‘foundation curriculum’). Even in the case where no amortization is possible, small model policies produce net savings in FLOPs over IID training, at the expense of marginal decreases in speedup in terms of learner FLOPs.
@@ 3.3 Unlocking compute-positive training
While model-based data-selection policies have demonstrated savings in terms of learner computation, none (to the best of our knowledge) produce overall savings when also accounting for the computation associated with the data selection process, which requires taking forward passes over candidate data. Although the cost of an inference pass $F$ is $\sim$1/3 the cost of a gradient update ($3F$) for most models, this cost scales with the proportion of data which is being rejected (e.g. retaining only 20% of the data requires 5 inference passes per trained batch). The requirements for compute-positivity can therefore be expressed as:
@@
where $F_{\text{act}}$ is the cost of scoring an example, $\rho$ is the number of examples scored per training example, and $\beta$ is the saving relative to IID training in terms of learner updates (see Appendix Section A.1 for more details). The RHS term is the cost of IID training per update. The first LHS term inside the brackets is the cost of the learner training during AL, the second term is the scoring cost and the last term outside the brackets the cost of training the reference model. We illustrate in Figure 2 the different contexts in which parts of this computation may be effectively amortized.
In the streaming and/or large-scale model training regime where data is not repeated nor seen before, in order to be compute-positive, either/all of the reference model training, actor scoring and learner efficiency $\beta$ terms must be made smaller to produce net savings vs. IID. Typical prioritization schemes can produce saving on the order of 50% (i.e. $\beta=0.5$), suggesting that savings must also be made by down-scaling the other terms $F_{\text{act}}$ and $F_{\text{ref}}$.
Cost of easy-reference scoring. While both the cost of the reference model and example scoring can be scaled down in the case of easy-reference scoring ($F_{\text{act}}=F_{\text{ref}}$, see equation 3), it is unclear whether the efficiency gains $\beta$ are robust to this down-scaling (see section 4.2).
Cost of RHO learnability scoring. The original definition of learnability scores [28] requires inference passes through both the learner and a reference model ($F_{\text{act}}=F_{\text{ref}}+F_{\text{learn}}$, see equation 4), meaning that although the cost of the reference model can be scaled down, the cost of example scoring cannot.
Cost of ClassAct / ActiveCLIP. For this reason, we explore whether replacing the learner model in term 1 of equation 4 with a much smaller model can still produce comparable learner efficiency gains to those already observed. Specifically, we introduce a third “online” model, which has the same architecture and size as the reference model, but is trained in parallel with the learner (see Algorithm 1). In this case, the cost of scoring examples reduces to:
@@
and can be arbitrarily scaled down along with the reference model. We instantiate our method for two canonical pre-training tasks: visual classification and multimodal learning, which we call ClassAct and ActiveCLIP respectively.
@@
3:    $X\sim\mathcal{D}$, where $|X|=B$triangleright Sample IID
@@ 3.4 Losses for canonical visual pre-training tasks
For visual classification, we use the standard cross-entropy loss for both actors and learners, i.e.
$\ell_{\textrm{act}}({\bm{x}}_{i},{\bm{y}}_{i}|\theta)=\ell_{\textrm{learn}}({\bm{x}}_{i},{\bm{y}}_{i}|\theta)=\textrm{XEnt}({\bm{y}}_{i},{\bm{l}}_{i}(\theta)),$ (8)
where ${\bm{l}}_{i}(\theta)$ are the logits produced by the model. For multimodal learning, learners optimize the contrastive loss $\ell_{\textrm{learn}}=\ell^{\textrm{im,txt}}_{\textrm{learn}}+\ell^{\textrm{txt,im}}_{\textrm{learn}}$, whereas the actor loss $\ell_{\textrm{act}}$ is simply the dot-product similarity between image and text embeddings:
@@
where ${\bm{z}}^{\textrm{im}}_{i}=f^{\textrm{im}}({\bm{x}}_{i};\theta)$ and ${\bm{z}}^{\textrm{txt}}_{i}=f^{\textrm{txt}}({\bm{x}}_{i};\theta)$ are image and text embeddings respectively, and $\ell^{\textrm{txt,im}}_{\textrm{learn}}({\bm{x}}_{i};\theta)$ is defined analogously to $\ell^{\textrm{im,txt}}_{\textrm{learn}}({\bm{x}}_{i};\theta)$.

### v4

@@ Abstract
Power-law scaling indicates that large-scale training with uniform sampling is prohibitively slow. Active learning methods aim to increase data efficiency by prioritizing learning on the most relevant examples. Despite their appeal, these methods have yet to be widely adopted since no one algorithm has been shown to a) generalize across models and tasks b) scale to large datasets and c) yield overall FLOP savings when accounting for the overhead of data selection. In this work we propose a method which satisfies these three properties, leveraging small, cheap proxy models to estimate “learnability” scores for datapoints, which are used to prioritize data for training much larger models. As a result, models trained using our methods – ClassAct and ActiveCLIP – require 46% and 51% fewer training updates and up to 25% less total computation to reach the same performance as uniformly-trained visual classifiers on JFT and multimodal models on ALIGN, respectively. Finally, we find our data-prioritization scheme to be complementary with recent data-curation and learning objectives, yielding a new state-of-the-art in several multimodal transfer tasks.
@@ 1 Introduction
Power-law scaling for vision and language models [19, 43] indicates that incremental improvements in model performance require orders of magnitude increases in computation. One of the key features of these empirical power-laws is that training data is sampled uniformly. In contrast, active data selection prioritizes computation on the data that maximally contributes to task performance [21, 24, 34], with the ultimate goals of improving data efficiency and reducing the cost of training. However, active data selection has yet to become a mainstay of large model training, since no existing algorithm satisfies a) robustness to the choice of model and training task, b) scalability to large datasets and architectures, and c) favourable end-to-end compute efficiency vs. training with uniform samples.
@@
Figure 1: Active learning accelerates large-scale visual understanding. For large-scale classification and multimodal learning tasks, prioritised training on data selected using our active selection methods ClassAct (left) and ActiveCLIP (right) requires significantly fewer updates to reach the final performance of uniform training.
@@
In the first instance, data selection based on hand-engineered filters (e.g. removing incorrectly shaped images or that only contain a single colour [2]) can trivially improve training efficiency at minimal computational overhead. However, such heuristics are limited in their effectiveness by the expertise of the human designer, and are not guaranteed to transfer to training of different models, data modalities, or tasks and incur significant effort to develop and tune.
@@
In contrast, model-based curation methods, which use the loss of the model itself to score examples, have shown promise by focusing training on ‘hard’ and omitting ‘easy’ data [35], but also the opposite (to exclude noise or other low-quality data) in both language modeling [13] and multimodal learning [15]. However, these methods often spend as much computation on the curation of datasets as is gained from subsequent pretraining, making them less compute-efficient than training on uniformly-sampled data. Finally, while several compute-efficient methods have been successfully deployed at small scale [8], they generally do not scale to even medium-sized datasets such as ImageNet.
In this work, we propose an algorithm that satisfies the three properties of generality, scalability, and compute-positivity. The proposed framework uses small proxy models to compute learnability scores for candidate training data, resulting in significant training efficiency gains at an almost negligible overhead over standard uniform-training. We test two instantiations of this framework, ClassAct and ActiveCLIP / ActiveSigLIP, for large scale classification and multimodal pretraining, respectively. Our findings are summarized below:
@@
Benchmarking heuristics for large-scale pretraining: We investigate loss- and learnability-based prioritization [15, 35, 27] for large-scale classification and find that pretrained reference models are an essential component for accelerating learning, producing efficiency gains of up to 46%.
Generalizing selection policies across model scale: Secondly, we show that smaller models act as effective proxies for much larger ($\sim$1000$\times$) models in the context of learnability but not loss-based scoring, resulting in compute-positive gains of up to 25% over uniform training, a first in the context of large-scale pretraining.
@@
Accelerating multimodal pretraining: Using reference models pretrained on small, clean datasets, we substantially accelerate pretraining on much larger, noisier datasets. Moreover, we find ActiveCLIP to be complementary to recent data-curation techniques [12] and learning objectives [44], yielding a new state-of-the-art in several multimodal understanding tasks.
@@
Amortizing data selection policies: Data-selection policies trained on one task can also accelerate the training of subsequent models on different but related tasks, suggesting that such policies can be easily derived from pre-trained models.
@@
Simplifying active-learning with online reference models: Lastly, we demonstrate that pre-trained reference models may not be necessary at all, where these models are small and can be trained in parallel on larger batches than the learner model, while remaining compute-positive.
@@ 2 Related Work
Data pruning. One approach to data-selection is to identify and sub-select data ahead of training. For example, [28] and [35] show that the training loss and gradients can be used to discard large portions of small-to-medium sized datasets (e.g. CIFAR10 and ImageNet) with little loss in performance. These methods have since been deployed for the curation of web-scale datasets in both language modeling [26] and multimodal learning [12, 1, 25], demonstrating large reductions in the amount of data required together with performance improvements. However, in the single-epoch regime that is becoming typical of large model training [16, 19], pre-filtering can be as expensive as learning from it, a shortcoming which we address in this work. Nevertheless, we show that our method for dynamic data selection is complementary to and benefits from such data curation techniques.
Online active learning. Unlike data-pruning, online active learning continuously filters data throughout training and applies naturally to the semi-infinite, single-epoch regime. Online Batch Selection [22] scores and filters using the learner model, which has the theoretical advantage that the importance of data can be determined relative to the current state of the learner. In terms of metrics, the Reducible Holdout Loss (RHO) [27] also uses the concept of a reference model to identify learnable data points not yet well represented. Other proposed heuristics include memorization for long-tailed data [11] and assigning “complexity” scores based on the number of times the example is forgotten during training [38]. None of these approaches however have demonstrated that the cost of scoring can be reduced to the point of justifying learner efficiency gains.
Compute-efficient data selection. Several works have demonstrated the benefits of selecting data based on simple heuristics, such as low-level image properties [2] or proximity to high-quality text corpora [29, 3, 7, 40]. While cheap to compute, these statistics often require domain-specific knowledge which limits their applicability across tasks. Domain-agnostic methods such as core-sets alleviate this by selecting data based on the geometry of their embeddings [14, 4] which can be efficiently computed, however these algorithms generally do not scale to large-scale datasets [8]. Most related to our work is DoReMi [39] which uses domain-general, scalable, and compute-efficient proxy models for the simpler problem of determining optimal data-mixtures for the subsequent training of a larger language model.
@@ 3.1 Data selection as prioritized replay
We use online batch selection [22] to apply our scoring heuristics to standard visual learning tasks: firstly, we sample uniformly from the training set ${\bm{x}}_{i}\stackrel{{\scriptstyle\mathcal{U}}}{{\sim}}\mathcal{D}$ and compute a score $s_{i}=s({\bm{x}}_{i}|\theta)\in\mathbb{R}$ to each data point ${\bm{x}}_{i}$ using model parameters $\theta$. Given a large enough collection of scored examples stored in a memory bank =$\mathcal{M}=\{{\bm{x}}_{i}\}_{i\in(0,\dots,M-1)}$, we then sample non-uniformly according to their scores ${\bm{x}}_{i}\stackrel{{\scriptstyle\pi}}{{\sim}}\mathcal{M}$ [31], where $\pi({\bm{x}}_{i})=\text{Softmax}(\{s_{i}\}_{i\in(0,...,M-1)})$. A batch of such examples is used to update the learner model. Following convention in reinforcement learning, we refer to the scoring and target models as actors and learners respectively.
@@ 3.2 Statistics for data selection
can use the current parameters of the learner $\theta^{t}$ or those of a fixed model $\theta^{*}$. The opposite argument can been made for favoring examples that are easily solved by a well-trained model, as such a prioritization removes the noisy examples present in large-scale datasets:
@@
This scheme is commonly used in multimodal learning for identifying high-quality examples with pre-trained models [15, 33, 32].
@@
A special case of learnability scores (the RHO loss, [27]) uses a model $\theta^{\textrm{ho}}$ specifically trained on a held-out dataset to ensure the independence of its predictions from those of the current learner $s^{\textrm{learn}}({\bm{x}}_{i}|\theta^{t},\theta^{\textrm{ho}})$. We assess in Section 4.2 whether this is necessary when training on large-scale image datasets.
Figure 2: Amortizing the cost of data selection. Drawn to scale: length of bars indicates number of FLOPs required to reach the accuracy of a ViT-L trained with uniform sampling (“ViT-L Uniform Sampling”, see Figure 4). Expensive model policies (e.g. a ViT-B scores data for the ViT-L learner, or ‘B $\rightarrow$ L’) produce large learner speedups, at the expense of the additional FLOPs associate with data selection. This overhead can be reduced by deriving the data-selection policies from smaller models (e.g. ViT-S, ViT-Ti or ViT-Mu score data for the ViT-L learner), at the expense of marginal decreases in the learner speedup. Costs could be additionally amortized by using off-the-shelf reference models, removing the need to train from scratch (yellow). Since the reference model is fixed throughout training, scores can be assigned once to a ‘foundation dataset’ and amortized across many training runs (lime green; [35, 27]). Since the online model is independent of the learner model and generalizes across scale, data selection policies can also be distilled as a fixed ordering of a given dataset (a ‘foundation curriculum’).
@@ 3.3 Unlocking compute-positive training
Scoring requires inference passes over the actor and learner models. We assume that gradient updates to cost 3x inference passes. The cost of scoring data $F$ thus scales with the proportion of data which is being rejected (e.g. retaining only 20% of the data requires 5 inference passes per trained batch). The requirements for compute-positivity can therefore be expressed as:
@@
where $F_{\text{act}}$ is the cost of scoring an example, $\rho$ is the number of examples scored per training example, and $\beta$ is the efficiency gain (in terms of the learner update) relative to training with uniform sampling (see Appendix Section 0.A.1 for more details). The RHS term is the cost per update of uniform sampling. The first LHS term inside the brackets is the cost of the learner training during AL, the second term is the scoring cost and the last term outside the brackets the cost of training the reference model. We illustrate in Figure 2 the different contexts in which parts of this computation may be effectively amortized.
In the large-scale training regime where data is neither repeated nor seen before, compute-positivity requires that either or all of the reference model training, actor scoring and learner efficiency $\beta$ terms must be made smaller to produce net savings relative to uniform sampling. Typical prioritization schemes can produce saving on the order of 50% (ie $\beta=0.5$), suggesting that savings must also be made by down-scaling the other terms $F_{\text{act}}$ and $F_{\text{ref}}$.
Cost of easy-reference scoring. While both the cost of the reference model and example scoring can be scaled down in the case of easy-reference scoring with small models ($F_{\text{act}}=F_{\text{ref}}$, see Equation 2), it is unclear whether the efficiency gains $\beta$ are robust to this down-scaling (see section 4.2).
Cost of RHO learnability scoring. The original definition of learnability scores [27] requires inference passes through both the learner and a reference model ($F_{\text{act}}=F_{\text{ref}}+F_{\text{learn}}$, see Equation 3), meaning that although the cost of the reference model can be reduce by using a smaller model, the cost of example scoring cannot.
Cost of ClassAct / ActiveCLIP. For this reason, we explore whether replacing the learner model in term 1 of Equation 3 with a much smaller model can still produce comparable learner efficiency gains to those already observed (see Appendix Algorithm 1). Specifically, we introduce a third “online” model, which has the same architecture and size as the reference model, but is trained in parallel with the learner. In this case, the cost of scoring examples reduces to:
@@
and can be scaled down along with the reference model. We instantiate our method for two canonical pre-training tasks: visual classification and multimodal learning, which we call ClassAct, ActiveCLIP and ActiveSigLIP respectively. In Appendix Section 0.A.6, we describe an asynchronous active learning framework where scoring and learning is performed in parallel on separate devices. This setup does not affect the FLOP calculations but can mitigate the overhead in time even when using non-approximate scoring.
@@
3:    $X\sim\mathcal{D}$, where $|X|=B$triangleright Sample uniformly
@@ 3.4 Losses for canonical visual pre-training tasks
For visual classification with ClassACT, we use the standard cross-entropy loss for both actors and learners. For multimodal learning with ActiveCLIP, learners optimize the contrastive loss $\ell_{\textrm{learn}}=\ell^{\textrm{im,txt}}_{\textrm{learn}}+\ell^{\textrm{txt,im}}_{\textrm{learn}}$, whereas the actor loss $\ell_{\textrm{act}}$ is simply the dot-product similarity between image and text embeddings:
@@
where ${\bm{z}}^{\textrm{im}}_{i}=f^{\textrm{im}}({\bm{x}}_{i};\theta)$ and ${\bm{z}}^{\textrm{txt}}_{i}=f^{\textrm{txt}}({\bm{x}}_{i};\theta)$ are image and text embeddings respectively, and $\ell^{\textrm{txt,im}}_{\textrm{learn}}({\bm{x}}_{i};\theta)$ is defined analogously to $\ell^{\textrm{im,txt}}_{\textrm{learn}}({\bm{x}}_{i};\theta)$. Similarly, ActiveSigLIP instead uses the sigmoid loss [44] for the learner’s objective and $\ell_{\textrm{act}}$ for scoring.